OpenStudy (anonymous):

change (-5,-7) into polar form and round both to the nearest tenth

OpenStudy (hba):

We have (x,y)=(-5,-7) We have to change it to (r,θ)

OpenStudy (hba):

To calculate use, \[\huge\ r = \sqrt{x² + y²}\] \[\huge\ θ = \arctan(y/x)\]

OpenStudy (anonymous):

I'm a little confused. Can you walk me through it?

OpenStudy (anonymous):

r^2=(-5)^2+(-7)^2 r^2=25+49 r^2=74 r=sqrt 74 but you must use a calculator to find θ.

OpenStudy (hba):

@marsss I guess i explained you something back there :/

OpenStudy (anonymous):

but she wanted to walk through :)

OpenStudy (hba):

@marsss -_-

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

welcome :)